# Spiral states, first-order transitions and specific heat multipeak phenomenon in $J_{1}$-$J_{2}$-$J_{3}$ model: A Wang-Landau algorithm study

### v1

@@ Abstract
The classical $J_{1}$-$J_{2}$-$J_{3}$ Ising model on the honeycomb lattice is important for understanding frustrated magnetic phenomena in materials such as $FePS_{3}$ and $Ba_{2}CoTeO_{6}$, where diverse phases (e.g., striped, zigzag, armchair) and magnetization plateaus have been experimentally observed. To explain the experimental results, previous mean-field studies have explored its thermal phase transitions, identifying armchair phases and striped phases, but their limitations call for more reliable numerical investigations. In this work, we systematically revisit the classical $J_{1}$-$J_{2}$-$J_{3}$ Ising model using the Wang-Landau algorithm. We find that the armchair (AC) phase, previously reported in mean-field and experimental studies, actually coexists with the spiral (SP) phase, with their combined degeneracy reaching 20-fold (4-fold for the AC states and 16-fold for the spiral states). The phase transitions and critical exponents are studied at different interaction values. We observe first-order phase transitions, continuous phase transitions, and even the multipeak phenomenon, i.e., Schottky-like specific-heat anomalies in frustrated systems. These results clarify the nature of phases and phase transitions in frustrated Ising systems and their exponents, and additionally provide inspiration for experimental efforts to search for the spiral state and Schottky-like anomalies.
@@ I Introduction
The main findings are as follows. The AC phase revealed in MF studies [13, 14] and experiments [11] coexists with the spiral (SP) phase in the intermediate parameter regime: the AC state is 4-fold degenerate, while the spiral state is 16-fold degenerate, leading to a total 20-fold degeneracy. Another key result is that, guided by the MF phase diagram, we accurately and reliably determine the ST-PM and AC+SP-PM phase transition boundaries at $J_{3}/J_{1}=1.2,0.6,0.2$, identify their transition types, and derive the corresponding critical exponents via the Wang-Landau algorithm. Additionally, in the strong frustration regime $J_{2}/J_{1}=-1.5$ and $J_{3}/J_{1}=0.2$ the specific heat exhibits a multi-peak behavior [22], which is related to the coexistence of the AC and SP phases.
@@ III.2 The phase diagram at $T>0$
Figure 11: At $J_{3}/J_{1}=0.2$, the temperature dependences of (a) the energy, (b) the specific heat, as well as (c) the double-peak distribution of the energy. Left column: $J_{2}/J_{1}=-1$; Right column: $J_{2}/J_{1}=-1.5$.
@@ III.2.4 The multipeak phenomenon of specific heat
As described in Ref. [22], the commonly observed multiple-peak behavior of specific heat corresponds to the Schottky anomaly [68, 69, 70], where the specific heat exhibits an anomalous sharp peak in addition to the conventional one. This phenomenon exists in non-frustrated systems, such as the classical spin-1 system, e.g., the Blume-Capel model [44]. In real solid materials, the spin characteristics of the Blume-Capel model show a correspondence with the actual Material. Taking the Fe-Mn-Al ternary alloy as an example [71], the Fe-Fe interactions in this alloy are ferromagnetic; the Mn-Mn and Mn-Fe interactions are antiferromagnetic; Al acts as a magnetic dilution element with a spin value of 0. For frustrated real materials, the multi-peak Schottky effect has been experimentally observed [72, 73, 74, 75, 76].
From the perspective of phenomenological theory, Ref. [22] states that a system with $n$ peaks should possess $n$ parameters. For the system with size 24 investigated in this paper, three peaks are indeed observed, which correspond to the three parameters $J_{1}$, $J_{2}$, and $J_{3}$.
@@ IV Conclusion and discussion
Regarding phase transitions, our WL results reveal distinct behaviors across parameter regimes: for $J_{3}/J_{1}=1.2$, the stripe (ST)-paramagnetic (PM) transition hosts a tricritical point near $J_{2}/J_{1}\approx-0.75$, while the determination of more precise values requires extensive calculations and analyses [77, 78, 79, 80, 81, 82, 83], which fall beyond the scope of this paper. For $J_{3}/J_{1}=0.6$, the ST-PM transition is either first-order or weak first-order; when $J_{3}/J_{1}=0.2$, the AC+SP-PM transition also follows first-order behavior. In the strong frustration regime ($J_{2}/J_{1}=-1.5$), we further observe multi-peak behavior in the specific heat. When $J_{3}=0$, our findings align with previous Metropolis and parallel tempering studies of the $J_{1}$-$J_{2}$ honeycomb model [66], confirming the consistency of our approach.
@@ Appendix A A snapshot in the real honeycomb lattices
Acknowledgements: W. Z. was supported by the Hefei National Research Center for Physical Sciences at the Microscale (Grant No. KF2021002) and the Fundamental Research Program of Shanxi Province (Grant Nos. 202303021221029). Y. J. was supported by the National Natural Science Foundation of China (Grant No. 12275263) and the Natural Science Foundation of Fujian Province of China (Grant No. 2023J02032).

### v2

# Spiral states, first-order transitions and specific heat multipeak phenomenon in $J_{1}$-$J_{2}$-$J_{3}$ Ising model: A Wang-Landau algorithm study

@@ Abstract
The classical $J_{1}$-$J_{2}$-$J_{3}$ Ising model on the honeycomb lattice is important for understanding frustrated magnetic phenomena in materials such as FePS3 and Ba2CoTeO6, where diverse phases (e.g., striped, zigzag, armchair) and magnetization plateaus have been experimentally observed. To explain the experimental results, previous mean-field studies have explored its thermal phase transitions, identifying armchair phases and striped phases, but their limitations call for more reliable numerical investigations. In this work, we systematically revisit the classical $J_{1}$-$J_{2}$-$J_{3}$ Ising model using the Wang-Landau algorithm. We find that the armchair (AC) phase, previously reported in mean-field and experimental studies, actually coexists with the spiral (SP) phase, with their combined degeneracy reaching 20-fold (4-fold for the AC states and 16-fold for the spiral states). The phase transitions and critical exponents are studied at different interaction values. We observe first-order phase transitions, continuous phase transitions, and even the multipeak phenomenon in frustrated systems. These results clarify the nature of phases and phase transitions in frustrated Ising systems and their exponents, and additionally provide inspiration for experimental efforts to search for the spiral state and specific-heat multipeak phenomenon.
@@ I Introduction
The main findings are as follows. The AC phase revealed in MF studies [13, 14] and experiments [11] coexists with the spiral (SP) phase in the intermediate parameter regime: the AC state is 4-fold degenerate, while the spiral state is 16-fold degenerate, leading to a total 20-fold degeneracy. Another key result is that, guided by the MF phase diagram, we accurately and reliably determine the ST-PM and AC+SP-PM phase transition boundaries at $J_{3}/J_{1}=1.2,0.6,0.2$, identify their transition types, and derive the corresponding critical exponents via the Wang-Landau algorithm. Additionally, in the strong frustration regime $J_{2}/J_{1}=-1.5$ and $J_{3}/J_{1}=0.2$ the specific heat exhibits a multi-peak behavior [22].
@@ III.2 The phase diagram at $T>0$
Figure 11: At $J_{3}/J_{1}=0.2$, the temperature dependences of (a) the energy, (b) the specific heat, as well as (c) the double-peak distribution of the energy for various system sizes. Left column: $J_{2}/J_{1}=-1$; Right column: $J_{2}/J_{1}=-1.5$.
@@ III.2.4 The multipeak phenomenon of specific heat
As described in Ref. [22], the commonly observed multiple-peak behavior of specific heat corresponds to the Schottky anomaly [68, 69, 70], where the specific heat exhibits an anomalous sharp peak in addition to the conventional one. This phenomenon exists in non-frustrated systems, such as the classical spin-1 system, e.g., the Blume-Capel model [44].For frustrated real materials, the multi-peak Schottky effect has been experimentally observed [71, 72, 73, 74, 75]. From the perspective of phenomenological theory, Ref. [22] states that a system with $n$ peaks should possess $n$ parameters. For the system with size 24 investigated in this paper, three peaks are indeed observed, which correspond to the three parameters $J_{1}$, $J_{2}$, and $J_{3}$.
However, the multiple peaks in the specific heat here are very likely a finite-size effect. Such effect has been studied in the $J_{1}-J_{2}-J_{3}$ Ising model in the square lattices by the MC method [76]. Subsequently, both the axial next-nearest-neighbor Ising (ANNNI) model and the biaxial next-nearest-neighbor Ising (BNNNI) model exhibit multiple peaks in their specific heat. In particular, for the ANNNI model with $J_{2}=0$ and finite $J_{3}$ along the horizontal axis, and the BNNNI model with $J_{2}=0$ and finite $J_{3}$ along both axes, a multi-peak structure in the specific heat is observed (see Fig. 1 in [77] and Fig. 10 in [78]). This feature is shown to be a finite-size signature characteristic of the incommensurate phase, evidenced by: (i) the number of peaks increases monotonically with increasing system size; and (ii) finite-size scaling analysis reveals that the extrapolated peak temperatures converge to a single point [79]. Due to the substantial computational demands in these strongly frustrated parameter regimes, our calculations are currently limited to $L=24$. The investigation of specific heat peaks for larger system sizes is therefore left to future work.
@@ IV Conclusion and discussion
Regarding phase transitions, our WL results reveal distinct behaviors across parameter regimes: for $J_{3}/J_{1}=1.2$, the stripe (ST)-paramagnetic (PM) transition hosts a tricritical point near $J_{2}/J_{1}\approx-0.75$, while the determination of more precise values requires extensive calculations and analyses [80, 81, 82, 83, 84, 85, 86], which fall beyond the scope of this paper. For $J_{3}/J_{1}=0.6$, the ST-PM transition is either first-order or first-order; when $J_{3}/J_{1}=0.2$, the AC+SP-PM transition also follows first-order behavior. In the strong frustration regime ($J_{2}/J_{1}=-1.5$), we further observe multi-peak behavior in the specific heat [76, 77]. When $J_{3}=0$, our findings align with previous Metropolis and parallel tempering studies of the $J_{1}$-$J_{2}$ honeycomb model [66], confirming the consistency of our approach.
@@
Acknowledgements: We thank Yi Hu and Patrick Charbonneau for helpful discussions. W. Z. was supported by the Hefei National Research Center for Physical Sciences at the Microscale (Grant No. KF2021002) and the Fundamental Research Program of Shanxi Province (Grant Nos. 202303021221029). Y. J. was supported by the National Natural Science Foundation of China (Grant No. 12275263) and the Natural Science Foundation of Fujian Province of China (Grant No. 2023J02032).